Of the height of a rectangale is 9 5/6 mm. and the base is 1 1/2 mm, what is the perimeter of the rectangle?
Answer by jorel1380(3719)

2(9 5/6)+2(1 1/2)=P
118/6+3=P
118/6+18/6=P
P=136/6 mm..